What is the slope y-intercept x-intercept of 5x-3y=5,x=7y-9, 2x-5y=1
Answer by Alan3354(69443)

What is the slope y-intercept x-intercept of 5x-3y=5,x=7y-9, 2x-5y=1
-----
5x-3y=5
Solve for y.
y = (5/3)x + b
The slope is the coefficient of x, 5/3
b is the y-intercept.
----
5x-3y=5
At the y-int, x=0
5*0-3y = 5
y = -5/3 or (0,-5/3) is the y-int.
---
At the x-int, y=0
5x-3*0 = 5
x = -5 or (-5,0) is the x-int.
======================
Do the other 2 the same way.